### v1

# Multi-Microphone Speech Emotion Recognition using the Hierarchical Token-semantic Audio Transformer Architecture

Abstract
Most emotion recognition systems fail in real-life situations (“in the wild” scenarios) where the audio is contaminated by reverberation. Our study explores new methods to alleviate the performance degradation of Speech Emotion Recognition (SER) algorithms and develop a more robust system for adverse conditions. We propose processing multi-microphone signals to address these challenges and improve emotion classification accuracy. We adopt a state-of-the-art transformer model, the Hierarchical Token-semantic Audio Transformer (HTS-AT), to handle multi-channel audio inputs. We evaluate two strategies: averaging mel-spectrograms across channels and summing patch-embedded representations. Our multi-microphone model achieves superior performance compared to single-channel baselines when tested on real-world reverberant environments.
keywords:
speech emotion recognition, human-robot interaction
1 Introduction
Speech Emotion Recognition (SER) is widely studied in the literature. Most of the reported studies deal with clean speech data and do not consider additive noise and reverberant environments typical to real-life applications. Only a few studies address the influence of reverberation and noise on SER. In these studies, the reverberant data is artificially generated by convolving the clean utterances with Room Impulse Responsess, either simulated [1, 2] or recorded in a real environment with various reverberation levels [3, 4]. Noise may also be added to the reverberant signals.
A survey of Human-Robot Interaction (HRI) in a noisy environment can be found in [5, 6]. However, this survey does not address the SER task. Previous works have shown the significant challenges of detecting emotional speech in large and reverberant rooms. Reverberation can influence the speech signal and negatively affect the predicted results [7]. The literature on using multiple microphones for SER is very scarce. In [8], the robustness of Automatic Speech Recognition (ASR) systems to emotional speech in noisy conditions is addressed. A Binaural Emotional Speech Recognition (BESR) system is proposed, enabling the simultaneous acquisition of the speaker’s emotional state and transcribing the uttered speech signal.
Devices equipped with multi-microphones are widely available nowadays. Adding the spatial information may improve the performance of audio processing tasks, including SER. However, learning-based algorithms are challenged by differences between training and test conditions, specifically a change in the microphone array constellations, e.g., when the number of microphones in train and test conditions is different.
In early works in the SER domain, various architectures such as Convolutional Neural Networks (CNN), Deep Neural Networks (DNN), Recurrent Neural Networks (RNN), and Long Short-term Memory (LSTM) were employed. Subsequently, combinations of CNN and RNN layers emerged, showcasing enhanced performance as compared with traditional classification methodologies [9, 10, 11]. Additionally, the effectiveness of configurations utilizing blocks comprised of CNN, LSTM and Bidirectional Long Short-term Memory (BiLSTM), as elucidated in [12, 13], was proven effective. Most of the works mainly focus on unimodal learning of emotions, either text, speech, or video [14, 15, 16].
The Transformer architecture, initially formulated for Natural Language Processing (NLP) [17, 18], has found applications in the audio processing domain, including tasks such as speech separation [19] and audio classification [20]. The superior performance of the Audio Spectrogram Transformer (AST) model [20], an adaptation of the Vision Transformer (ViT) model [21], was demonstrated compared to CNN-based models. It is imperative to highlight that the AST model exclusively handles single-microphone data, whereas, in real-world scenarios, multiple microphones may often be available. A drawback of Transformer models is their reliance on large training data for convergence.
In the current work, we extend the HTS-AT architecture to accommodate multi-channel inputs, thus enhancing robustness against reverberations. Since only a limited amount of data is available for our task, we resort to fine-tuning already-trained models. Moreover, our scheme can be fine-tuned with a certain number of microphones and tested with another microphone constellation, including microphone positions and the number of microphones. We evaluated the SER performance using three datasets, Ryerson audio-visual database of emotional speech and song (RAVDESS) [22], Interactive Emotional Dyadic Motion Capture (IEMOCAP) [23], and Crowd-sourced Emotional Multimodal Actors Dataset (CREMA-D) [24]. We used real-life RIRs from the Acoustic Characterisation of Environments (ACE) Challenge [25] to add reverberation to the speech utterances.
2 Problem Formulation
Let $x(t)$ denote the anechoic signal in the discrete-time domain. An $M$-microphone array captures this signal after propagating in the acoustic enclosure. The received microphone signals are then given by $y_{i}(t)=\{x*h_{i}\}(t),i=1,2,\ldots,M$, where $h_{i}(t)$ is the RIR from the source position to the position of microphone $i$. The aim of this work is to classify the emotion given the observations $y_{i}(t);i=1,\ldots,M$.
3 Proposed Model
The proposed SER model is based on the Swin-Transformer [26],111github.com/microsoft/Swin-Transformer, a state-of-the-art hierarchical ViT [21] architecture, used for a variety of computer vision tasks, that utilizes shifted windows to capture long-range image dependencies. We adopted and modified a Swin Transformer architecture, namely the Hierarchical Token-semantic Audio Transformer (HTS-AT) [27], which aims to improve the performance and scalability of audio tasks, such as the AudioSet dataset [28]. The HTS-AT model is designed to achieve the best performance by reducing the number of parameters, requiring fewer GPU resources, and less training time than the AST architecture [27]. In the current work, we propose further modifying the HTS-AT architecture to better suit the audio processing requirements. The main change is to adapt the model for multi-channel audio processing, expanding its usefulness beyond the original single-channel design. Two alternative multi-channel pre-processing strategies are examined. The first strategy applies a summation of patch tokens derived from mel-spectrograms. The second strategy applies averaging of mel-spectrograms.
3.1 Preprocessing and Input Features
We assume that the sampling rate of the audio signals is 16 kHz. Each microphone signal is first analyzed by a short-time Fourier transform (STFT), with a window size of 1024 samples and a hop size of 160 samples. The STFT bins are then aggregated to construct the mel-spectrograms. After fusing the multi-channel information into a single stream, the regular HTS-AT architecture is applied.
3.2 Architecture
We will now elaborate on the HTS-AT architecture. The standard transformer architecture requires extensive computational resources due to the unmodified input token sequence length across all layers. This includes maintaining a large global self-attention matrix and calculating outputs and gradients at each step. The HTS-AT architecture is introduced to address these challenges. The HTS-AT is designed for supporting multiple audio tasks, e.g., classification, sound event detection, and source localization. It introduces two key architectural optimizations: a hierarchical structure and a windowed attention mechanism. The input audio mel-spectrogram is split into localized patch tokens using a convolutional Patch-Embed CNN layer of kernel dimensions $P\times P$, in which the patches are ordered by time segment and frequency bin. Then, tokens propagate through a series of Swin Transformer encoder groups. At the terminus of each group, a Patch Merging layer reshapes the token sequence into its original 2D mel-spectrogram. This layer merges neighboring patches and then embeds them back into a latent space of reduced length. Consequently, the memory requirements decrease exponentially with depth. Within each Swin Transformer block, attention is restricted to non-overlapping $M\times M$ squares, partitioning the token sequence. Calculating self-attention within each window subset substantially reduces computational complexity relative to full global attention while capturing localized relationships. As patch size increases downstream, windows encapsulate larger temporal and frequency contexts. Finally, the HTS-AT incorporates a token-semantic CNN layer after the last transformer block. This layer refines the output by grouping tokens, thus capturing information about their time frames and frequency bins. Consequently, this enhances classification by exploiting token relationships.
3.3 Multi-Channel Methods
We propose two multi-channel approaches that enable the HTS-AT model to handle single- and multi-channel inputs through fine-tuning. Both methods maintain HTS-AT’s core architecture to allow flexibility in microphone numbers during fine-tuning and evaluation.
Figure 1: Scheme of patch tokens mel-spectrogram segments.
Figure 2: Scheme of average mel-spectrogram.
Inspired by the study in [29], we embraced a Patch-Embed scheme where each channel undergoes a shared embedding layer followed by summation. This maintains flexibility in the microphone numbers used during fine-tuning and testing without altering the core HTS-AT architecture. Additionally, summing Patch-Embed outputs consolidates inter-channel information, enabling more robust representations suitable for distortion and reverberation conditions.
Patch-Embed Scheme: On the left side of Fig. 1, each of the $M$ channels is analyzed to generate mel-spectrograms, which are then concatenated along the channel depth. SpecAugment [30] is applied collectively across all $M$ mel-spectrograms. After this, each mel-spectrogram is reshaped to the dimensions of a $256\times 256$ image and passes through the shared Patch-Embed layer. Next, a summation operation is performed across all $M$-encoded channels, making this input suitable for the pre-trained HTS-AT. Finally, a single representation is derived, which is then sent to the feedforward mechanism of HTS-AT for emotion classification.
Average Mel-Spectrograms: The spectrogram averaging technique reduces reverberation effects by consolidating information from all channels. Since the model input is a mel-spectrogram image, averaging enhances frequency bias in the resulting image. This process happens at the start of the signal processing phase, before the Patch-Embed encoder layer. On the left side of Fig. 2, each channel is individually processed into a mel-spectrogram and then combined with other processed mel-spectrogram channels. Subsequently, averaging is performed based on the specified number of input channels, resulting in a transformation from a multi-channel to a single-channel representation. After this transformation, the mel-spectrogram is augmented and structured into an image. Following this, the image is sent to the Patch-Embed encoder layer, functioning as a feature tensor with dimensions of $4096\times 96$, and then progresses to the feedforward mechanism within the HTS-AT framework for emotion classification. It is worth noting that since the pre-trained model was specifically trained to handle a single channel, this characteristic enables the smooth integration of the mentioned representation into the model architecture. In contrast to the Patch-Embed Scheme approach, this method emphasizes preprocessing before patch embedding. As a result, it allows fine-tuning of a model with a predetermined number of microphones, making it easier to test with different microphone configurations. This feature provides flexibility and autonomy in the use of any number of microphones.
Table 1: The results of single-microphone fine-tuned HTS-AT model compared with the method in [12] on clean (non-reverberant) datasets. We present the weighted average accuracy (in percentage) of 20 models that achieved the highest accuracy on the validation set and tested on the test set.
Models RAVDESS IEMOCAP CREMA-D
Fine-tuned HTS-AT 90% 70.93% 75.86%
BiLSTM + Attention[12] 80% 66% -
4 Experimental Study
Datasets: Our study comprised three speech emotion recognition datasets, namely RAVDESS[22], IEMOCAP[23] and CREMA-D[24] datasets. The RAVDESS dataset comprises 24 actors, evenly distributed between male and female speakers, each uttering 60 English sentences. Hence, there are 1440 utterances expressing eight different emotions: ‘sad’, ‘happy’, ‘angry’, ‘calm’, ‘fearful’, ‘surprised’, ‘neutral’, and ‘disgust’. All utterances are transcribed in advance. Consequently, the emotions are more artificially expressed as compared to spontaneous conversation. In this dataset, we decided to union the emotions ‘neutral’ and ‘calm’ as representations of ‘neutral’. Therefore, we fine-tuned our model on seven classes instead of eight. A major drawback of the dataset is the small number of utterances. The IEMOCAP dataset comprises approximately 12 hours of speech and consists of conversations between two people that are either improvised or played according to a pre-determined transcript that was chosen to evoke different emotions. We fine-tuned our model on four classes: ‘happy’, ‘sad’, ‘angry’ and ‘neutral’. CREMA-D is a dataset of 7442 original clips from 91 actors comprising 48 male and 43 female actors. Speech utterances were selected from a set of 12 sentences. The sentences were presented using one of six different emotions ‘anger’, ‘disgust’, ‘fear’, ‘happy’, ‘neutral’ and ‘sad’.
For the multi-channel simulated RIR datasets, we first split the original three datasets into 80% of the data as a training set, 10% as a validation set, and 10% as a test set. Then, we used the gpuRIR Python code package222github.com/DavidDiazGuerra/gpuRIR to add reverberation and simulate multi-channel microphone signals (four channels) with reverberation time in the range of $T_{60}=200-800$ ms with different room sizes and randomized microphone locations. By doing so, we managed to enlarge the datasets. For RAVDESS, 6863 reverberant speech samples were generated for training, and 852 samples were used for validation. We follow the same procedure for IEMOCAP with 7356 train samples and 2107 validation samples and for CREMA-D with 5945 train samples and 1487 validation samples. Our objective encompasses evaluating our model in real-world reverberant environments. To achieve this, we conducted tests utilizing the ACE RIR database [25], which comprises seven distinct rooms characterized by varying dimensions and exhibiting diverse ranges of $T_{60}$ captured by a mobile phone equipped with three microphones. The model was fine-tuned with the synthesized RIRs and evaluated with the test sets of the various datasets convolved with the ACE RIRs.
Table 2: Results on the test sets of the RAVDESS, IEMOCAP, and CREMA-D datasets convolved with RIRs of three microphones from the ACE database. The ‘HTS-AT’ columns are fine-tuned on reverberant single-channel audio. The ‘Avg mel’ columns depict results where mel-spectrograms were averaged across four channels during fine-tuning and tested on three channels. The ‘Sum PE’ columns are the Patch-Embed Scheme approach fine-tuned and tested on three channels.
RAVDESS IEMOCAP CREMA-D
HTS-AT Avg mel Sum PE HTS-AT Avg mel Sum PE HTS-AT Avg mel Sum PE
Lecture Room 1 ($T_{60}=638$ ms) 77.3% 80.6% 81.3% 61.3% 67.0% 67.4% 63.2% 66.4% 67.4%
Lecture Room 2 ($T_{60}=1220$ ms) 77.3% 78.6% 82.0% 63.4% 66.1% 68.7% 65.4% 67.3% 66.0%
Lobby ($T_{60}=646$ ms) 78.6% 82.0% 84.0% 61.8% 64.0% 65.1% 64.2% 66.2% 66.8%
Meeting Room 1 ($T_{60}=437$ ms) 73.3% 80.6% 82.6% 62.9% 66.7% 68.7% 64.4% 65.7% 65.7%
Meeting Room 2 ($T_{60}=371$ ms) 82.0% 83.3% 85.3% 59.5% 66.3% 64.2% 65.6% 66.6% 66.8%
Office 1 ($T_{60}=332$ ms) 76.0% 83.3% 80.6% 63.6% 68.3% 67.2% 64.1% 68.8% 67.3%
Office 2 ($T_{60}=390$ ms) 78.6% 81.3% 82.6% 59.5% 64.7% 65.4% 62.6% 64.0% 64.1%
Algorithm Setup: The original HTS-AT model was pre-trained utilizing the AudioSet dataset, which comprises over two million audio samples. Each sample is 10 seconds long and is categorized into 527 distinct sound event categories. In order to adapt the pre-trained AduioSet model to RAVDESS, IEMOCAP, and CREMA-D, we adjusted the number of output classes to 7, 4, and 6, respectively. In the three datasets, pre-processing and warm-up strategy were carried out as demonstrated in [27] by providing the HTS-AT with 64 mel-bins to compute the STFT and mel-spectrograms features with 160 hop size and 1024 window size. We modified the original AdamW optimizer of HTS-AT to the traditional Adam optimizer with a learning rate of $1e^{-3}$, and batch size of 128. We set the number of epochs to 150 for all datasets. We used cross-entropy loss as the metric. We also follow the same original HTS-AT hyperparameter settings. The dimensions of the patch are set to $4\times 4$, the patch window length is 256 frames, and the attention window size is $8\times 8$. The architectural configuration comprises four network groups, each comprising 2, 2, 6, and 2 Swin-Transformer blocks, respectively. The initial patch embed is linearly projected to a dimension of $D=96$, and correspondingly, after each transformer group, the dimension increases exponentially to $7688D=$ 768, aligning seamlessly with the principles of AST. In part of our experiments, we reduced the depth of each network group by half, thereby reducing the number of trainable parameters by half to prevent overfitting. In addition, we added an early stopping strategy with a patience of 50 for RAVDESS and 25 for IEMOCAP and CREMA-D. The overall parameters for fine-tuning the datasets were 15.7M for the RAVDESS and 28.6M for IEMOCAP and CREMA-D. Since we performed fine-tuning, our models trained in less than two hours on one A6000 RTX GPU.
(a) IEMOCAP
(b) RAVDESS
Figure 3: Accuracy and Confidence Interval on test sets convolved with ACE RIR Lecture Room 2 ($T_{60}=1220~{}ms$). The results of two HTS-AT fine-tuned on clean and simulated RIRs datasets compared with [12] trained on clean datasets.
Results: We applied fine-tuning to the pre-trained HTS-AT AudioSet model using single-channel clean (non-reverberant) and multi-channel simulated RIRs datasets. Our starting point referred to the clean single-channel case, which will serve as one of the baselines. We also compared our results with another work that proposed a different architecture using BiLSTM together with the Attention mechanism [12]. Table 1 depicts the process of fine-tuning the model using the clean single-channel datasets. Furthermore, to complete establishing the baseline using the RAVDESS and IEMOCAP datasets, we fine-tuned another single-channel HTS-AT model by fine-tuning on the artificially reverberated datasets with a uniformly distributed $T_{60}=200-800$ ms. Then, we compared the three models by computing the Confidence Interval.333github.com/luferrer/ConfidenceIntervals The evaluation was performed on the reverberant test set using the ACE database with only one of the microphones. Examining Fig. 3 for both IEMOCAP RAVDESS datasets, it is evident that the two HTS-AT variants clearly outperform [12]. It is also clear that for RAVDESS, the HTS-AT fine-tuned with reverberant speech outperforms the HTS-AT fine-tuned using clean data. While this is also true for IEMOCAP, the differences are less significant.
We now turn to the evaluation of the multi-channel schemes. In all experiments, we tested all three datasets with the utterances convolved with the three RIRs of the ACE database. We only report the results for the remote source case to emphasize the reverberation effects. The average mel-spectrogram scheme was fine-tuned using four microphones. The Patch-Embed scheme was fine-tuned using three microphones. In both cases the RIRs were generated using the simulator, as explained above. All experimental procedures followed identical training-validation splits, maintaining consistent model configurations, sizes, and hyperparameters for each dataset.
Table LABEL:ACE_results provides a comparative assessment of three model variants, all fine-tuned with reverberant speech: 1) ‘HTS-AT’ - a single-channel scheme, 2) ‘Avg mel’ - mel-spectrograms averaged across four channels, and 3) ‘Sum PE’ - the Patch-Embed scheme. It is evident that the ‘Sum PE’ approach consistently outperforms the other two variants, although not by a large margin. Statistical analysis of the SER accuracy for four of the rooms is depicted in Fig. 4. Examining the accuracy bars and the confidence intervals confirms our findings regarding the consistent, albeit modest, improvements achieved by the ’Sum PE’ scheme. Importantly, these benefits come without a considerable increase in computational complexity compared to single-channel models.
(a) IEMOCAP ACE Lecture Room 1 ($T_{60}=638~{}ms$).
(b) RAVDESS ACE Meeting Room 1 ($T_{60}=437~{}ms$).
(c) CREMA-D ACE Office 1 ($T_{60}=332~{}ms$).
(d) IEMOCAP ACE Lecture Room 2 ($T_{60}=1220~{}ms$).
Figure 4: Accuracy and Confidence Interval on test sets convolved with four rooms from the ACE database. A slight improvement of the multi-channel schemes over the single-channel baseline is demonstrated, with a marginal advantage to the ‘Sum PE’ scheme.
5 Conclusions
In this paper, we presented a multi-microphone transformer-based model for SER in reverberant environments. Based on the HTS-AT architecture, the model employs two strategies for handling multi-channel audio inputs: averaging mel-spectrograms across channels and summing patch-embedded representations. When tested on real-world reverberant environments, the results show improved SER accuracy compared to single-channel schemes. By leveraging spatial information from multiple microphones, our model was able to exhibit a more robust behavior of the SER in challenging acoustic conditions. The consistent, but not large, improvements of the proposed multi-microphone schemes show promise for developing SER systems that can perform reliably in noisy and reverberant scenarios.